### v1

# From Calculation to Adjudication: Examining LLM judges on Mathematical Reasoning Tasks

Abstract
To reduce the need for human annotations, large language models (LLMs) have been proposed as judges of the quality of other candidate models. LLM judges are typically evaluated by measuring the correlation with human judgments on generation tasks such as summarization or machine translation. In contrast, we study LLM judges on mathematical reasoning tasks. These tasks require multi-step reasoning, and the correctness of their solutions is verifiable, enabling a more objective evaluation. We perform a detailed performance analysis and find that the used judges are mostly unable to improve task performance but are able to pick the better model. Our analysis uncovers a strong correlation between judgment performance and the candidate model task performance. We observe that judges tend to choose the model of higher quality even if its answer is incorrect. Further, we show that it is possible to use statistics, such as the task performances of the individual models, to predict judgment performance. In an ablation, we either swap or mask the candidate answers and observe that judges often keep the original judgment, providing evidence that judges incorporate writing style in their judgments. In summary, we find that regularities in the judgments are quantifiable using statistical measures and provide various angles on exploiting them. [^1] Code: git@github.com:AndSt/llm_judges.git
1 Introduction
The automatic evaluation of machine learning models promises to reduce the need for human annotations. Specifically, the LLM-as-a-judge paradigm Zheng et al. (2023) has gained traction, aiming to assess or compare the quality of generated texts automatically. This approach is beneficial for automated data labeling Tan et al. (2024), self-improvement of LLMs Wu et al. (2024), and ranking LLMs with respect to specific tasks Zheng et al. (2023).
Figure 1: In our problem setup two LLMs ($M_{A}$ and $M_{B}$), provide candidate answers for a math problem, and a judge LLM has to decide which one is correct. All three use chain-of-thought (CoT) reasoning Wei et al. (2022).
Much like judges in the real world, who are expected to be exact, fair, and unbiased, e.g., as defined in Bangalore Principles of judicial conduct Bangalore Principles, (2002), LLMs, when employed as judges, should be ethical and logical. Already the philosopher Aristotle argued that the virtuous actor exhibits the joint excellence of reason and character Kraut (2022). Previous works investigate properties and biases of LLM judges on generation tasks such as translation or summarization Kim et al. (2024b); Liu et al. (2024). These are typically evaluated using correlation with human annotators and are thus inherently subjective.
In this work, we investigate LLM judges on mathematical reasoning datasets (see Figure 1). These need complex multi-step reasoning, and the solution is verifiable, which allows us to investigate the relationship between judge and candidate models in a principled manner. We base our analysis on four large (more than 30B parameters) LLMs and four small (less than 10B) LLMs on three mathematical reasoning datasets.
Our experiments contain a detailed performance examination, confirming that larger models are generally better judges Zheng et al. (2023). We find that only the best-tested model, Qwen 2 72B, consistently improves task performance if we evaluate the judged samples, but all tested judges likely pick the better model for a given task.
We investigate subsets with one correct and one incorrect candidate answer. We uncover a correlation between judgment performance and task performance of the candidate models, showing that judges tend to select incorrect answers from better models. Thus, we hypothesize that judges have access and rely on the superior writing styles of larger models instead of solely analyzing the reasoning. When we divide the datasets into buckets of model agreement, we observe that agreement is a proxy for sample difficulty.
Motivated by these regularities, we analyze whether it is possible to predict judgment performance and find that task performances of judge and candidate LLMs explain most of the variance. We hypothesize that judges incorporate writing style into their judgments. Thus, we predict individual judgments using statistical and transformer-based models and achieve above-chance performance, supporting our hypothesis.
Lastly, we test how judgments are affected by perturbing numeric values in responses by 1) swapping results and 2) masking numeric values. Our findings reveal that judges largely retain original judgments, providing further evidence that judges, in large part, base their decisions on writing style. In summary, our contributions are as follows:
  1. 1.
    We conduct an in-depth performance analysis of LLM judges for mathematical reasoning tasks.
  2. 2.
    Our analysis reveals a correlation between the judgment and candidate task performance, providing a novel statistical angle on the analysis of LLM judges.
  3. 3.
    We show that statistics such as task performance or agreement of candidate models are indicative of judgment performance.
  4. 4.
    After systematically perturbing the candidate answers, we observe that judges often keep their original judgments, providing evidence that judgments are also based on writing style.
2 Related Work
2.1 LLM as Judges
Using LLMs as judges to evaluate text generated by LLMs, including their own outputs, has recently attracted significant interest because it reduces the need for human annotation Zheng et al. (2023). Commonly, large frontier models are used as judges. Applications include the automatic assessment of language model capabilities and, e.g., determining which model performs better on a given task Zheng et al. (2023) and reinforcement learning from AI feedback by automatically generating data for preference optimization Bai et al. (2022); Wu et al. (2024).
Various methods exist to make judgments Zheng et al. (2023); Liusie et al. (2024). One approach is pairwise selection Wang et al. (2024a), where two answers are presented, and the model is asked to select the better one. Another approach is pointwise grading Li et al. (2024), where the model is asked to assign a grade based on a predefined scale, and the answer with a better grade is chosen. Judgment prompts may involve reference solutions or not. Another body of research explicitly trains models to act as judges Kim et al. (2024a); Wang et al. (2024a) or closely related, as reward models Wang et al. (2024b); Li et al. (2024).
The effectiveness of LLMs as judges is typically assessed by measuring the correlation or overlap with human judgments Zheng et al. (2023); Kim et al. (2024b). In contrast, we focus on difficult tasks with a concrete final answer. Finally, we want to stress that several works caution for the use of LLM judges as experts Bavaresco et al. (2024); Koo et al. (2023); Raina et al. (2024). In a similar vein, we aim to understand regularities and their shortcomings.
2.2 Biases in LLM-as-a-judge
Human-annotated data inherently reflects the annotators’ biases and opinions. These biases can be detrimental or (intentionally) beneficial, depending on the goals of the annotation process Plank (2022). Similarly, several studies have explored the biases present in LLM judges:
One linguistic bias is ordering bias Zheng et al. (2023); Koo et al. (2023), where a judge gives a different answer depending on the order in which answers are presented. Panickssery et al. (2024) note that it is possible to interpret position bias as a sign that the model is unsure. There are multiple works Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024) that find evidence for self-bias or self-preference. Koo et al. (2023) provide a benchmark for analyzing cognitive biases. West et al. (2024) and Oh et al. (2024) explore the “Generative AI Paradox” where generating solutions is easier for the LLM than analyzing them, unlike humans who typically find analysis easier than generation.
In this work, we aim to establish a better understanding of underlying regularities that relate judgments to statistics such as model performance.
3 General Setup
In the following, we describe the problem setting, including the used notation, and the general experimental setting including used models and datasets.
3.1 Problem Description
In this work, we consider two models, denoted by $M_{A},M_{B}\in\mathbf{M}$, providing candidate solutions for a sample of a dataset $D$ and a judge model $M_{J}\in\mathbf{M}$, which is tasked to select, to “judge”, whether it prefers the solutions of the models $M_{A}$ or $M_{B}$. The solutions are represented by the random variables $A$ and $B$. We consider the events that solutions are true $(A=T)$, false $(A=F)$, or that their solution is the same $(A=B)$. We denote the judgment of the judge $M_{J}$ by the random variable $\Delta_{J}$, which can either be correct $(\Delta_{J}=T)$, incorrect $(\Delta_{J}=F)$ or choose a specific model $M_{A}$ $(\Delta_{J}=M_{A})$.
Given that the final answer is either correct or incorrect, we can break the probability of the judge making a correct judgment $P(\Delta_{J}=T|A,B,D)$ given a sample of a dataset $D$ and the answers of two models $M_{A},M_{B}$ down into the following four cases:
$P(\Delta_{J}=T|A,B,D)$ (1)
$=\sum_{(X,Y)\in C}P(\Delta_{J}=T|A=X,B=Y,D)P(A=X,B=Y|D)$
$=P(A=T,B=T|D)$
$+P(\Delta_{J}=T|A=T,B=F,D)P(A=T,B=F|D)$
$+P(\Delta_{J}=T|A=F,B=T,D)P(A=T,B=F|D)$
where $C=(T,F)^{2}$. Note that in cases where both answers are correct or incorrect imply that the judgment is also either correct or incorrect respectively, i.e., $P(\Delta_{J}=T|A=T,B=T)=1$ and $P(\Delta_{J}=T|A=F,B=F)=0$.
3.2 Datasets
The experiments encompass three mathematical reasoning datasets where models highly benefit from multi-step CoT reasoning. For all datasets, we use accuracy as the performance metric.
AQUA-RAT Ling et al. (2017) is a dataset to test the quantitative reasoning ability of LLMs. Unlike the other two datasets, the questions are multiple-choice. GSM8K Cobbe et al. (2021) consists of grade school math word problems. The answers are free-form numbers. MATH Hendrycks et al. (2021) contains challenging competition mathematics problems. Find more details in Appendix A.1
3.3 Models
We evaluate the performance of openly available LLMs, including four large models Qwen 2 72B Yang et al. (2024), Llama 3 70B AI@Meta (2024), Yi 1.5 34B Young et al. (2024), Mixtral 8x7B Jiang et al. (2024) and four small models, namely Llama 3 8B AI@Meta (2024), Gemma 1.1 7B Gemma Team et al. (2024), Mistral 7B v0.3 Jiang et al. (2023), and Mistral 7B v0.1 Jiang et al. (2023). We use the chat- or instruction-tuned model variants and test each model as a candidate answer generator and as a judge. More information is in Appendix A.2.
3.4 Inferences
This section describes the candidate answer generation and the judgment generation. Find more information on prompts and hardware details in Appendix A.
Candidate answer generation.
To judge two candidate answers (including of the same model), we sample two initial CoT solutions for each model using 4-shot prompting. We set the temperature to $0.9$ to get two different solutions.
Qwen 2 72B Llama 3 70B Yi 1.5 34B Mixtral 8x7B Llama 3 8B Gemma 1.1 7B Mistral 7B v0.3 Mistral 7B v0.1
(1) $P(\Delta_{J}=T|A,B,D)$ AQUA_RAT 66.05 55.98 62.2 57.08 51.59 53.52 54.96 51.41
GSM8K 77.06 72.41 72.64 68.81 65.88 65.47 68.57 63.28
MATH 29.66 24.64 26.69 23.60 22.35 21.66 22.68 19.91
(2) $P(\Delta_{J}=T|A\neq B,D)$ AQUA_RAT 53.79 45.04 49.73 44.47 36.56 40.48 38.26 36.43
GSM8K 63.16 59.33 56.01 46.68 41.29 38.71 42.59 39.72
MATH 25.04 22.14 22.22 18.29 17.04 16.00 16.88 15.22
(3) $P(\Delta_{J}=T|\{A,B\}=\{T,F\},D)$ AQUA_RAT 73.13 64.63 68.66 63.26 52.74 58.04 54.54 52.45
GSM8K 85.65 81.32 76.84 64.97 57.87 54.16 59.58 55.48
MATH 80.48 73.01 71.37 61.03 58.04 54.70 56.27 50.91
Table 1: Performance of judge LLMs in three cases: (1) accuracy on all samples, (2) accuracy where models $M_{A}$ and $M_{B}$ disagree, and (3) accuracy where only one model is correct. Results are averaged over all pairs $(M_{A},M_{B})$, with the highest accuracy in bold and the second highest underlined.
Judgements.
We choose the first candidate generation for each model and generate judgments for all $36$ unique model combinations. [^2] We consider all pairs from the eight LLMs, including self-pairing, yielding $\binom{8+2-1}{2}=36$ combinations. If both models are the same, we take the second initial generation. We accommodate positional bias Zheng et al. (2023); Koo et al. (2023) by evaluating the two candidate answers in both possible orders for each question and then taking the average correctness of the judgments as the final assessment. The judge has to choose if the first or second answer is correct. The prompt is zero-shot and applies CoT, the temperature is set to 0 for deterministic generation results.
4 General Performance
The experiments have multiple degrees of freedom: judges, candidate models, and datasets. Therefore, we first examine judgments per dataset, and secondly, we investigate judgments per candidate model pair. Afterwards, we provide evaluations for two applied questions.
4.1 Performance per dataset
We begin by examining the judgment performance, i.e., how often the judge picks a correct answer, across different datasets. Therefore, we average the performance across all model pairs $(M_{A},M_{B})$.
Setup.
Table 1 considers three cases where each case focuses on a specific subset of the datasets: Case (1) investigates the observed task performance $P(\Delta_{J}=T|A,B,D)$ where we evaluate the task performance using the answers chosen by the judges. Note that this includes samples where both candidate models give the same answer. Case (2) asks how often judges choose a correct answer given that the answers differ, i.e., $P(\Delta_{J}=T|A\neq B,D)$. Note that this may (and often does) include cases where both answers are incorrect. Case (3) gives the probability that the judge chooses the correct answer given that one answer is correct, and the other answer is incorrect, formally $P(\Delta_{J}=T|A\neq B,T\in\{A,B\},D)$.
Results.
We observe that large models outperform smaller models. Specifically, we see that Qwen 2 72B is the best judge, followed by Yi 1.5 34B. The performance of Llama 3 70B is, on average, comparable to that of Yi 1.5 34B. Note that performance in Case (1) and Case (2) is often quite low, especially for MATH, as there are many cases where the judge can only choose wrong answers. Importantly, we observe that smaller models with fewer than 10B parameters are unreliable judges. Especially, in Case (3), where a correct answer is provided, smaller models only achieve an accuracy of around 55%, barely better than random chance. Therefore, we focus on the four larger models as judges in the subsequent analysis.
(a) Qwen 2 70B
(b) LLama3 70B
(c) Yi 1.5 34B
(d) Mixtral 8x7B
Figure 2: Observed performance $P(\Delta_{J}=T|A,B,D)$ of four judge LLMs (a-d) in evaluating various model pairs, averaged across all datasets.
4.2 Performance per model combination
The comparative performance of model pairs offers insights into which model is better for the specific task or which combination of models yields the best results.
Setup.
Figure 2 illustrates the final performance $P(\Delta_{J}=T|A,B)$, indicating the probability of a judge choosing a correct answer given two models $A$ and $B$. The results are averaged over datasets and presented as an upper triangular matrix due to symmetry. If both models in a pair are the same, $A=B$, we employ the second response generated with temperature sampling to introduce variation. We report the performance of all models used as judges in the Appendix B in Table D.
Results.
We observe that the best performance is achieved when both the candidate answers and the judge are the highest-performing model, Qwen 2 72B. An analysis of the first rows (cf. Figure 2) reveals a notable trend: The final performance declines when comparing the output of a strong model against a mediocre model (e.g., Llama3-8B) but then improves again when compared against the weakest model. This suggests that judging becomes more challenging when distinguishing between the correct answers of a strong model and the incorrect answers of a mediocre one, compared to discerning the outputs of a bad one.
4.3 Do judges elicit task improvement?
Figure 3: Amount of model pairs $(M_{A},M_{B})$ where the answers chosen by the judge achieve a higher task performance than the models individually (green). The blue bar only considers models where the judge is at least as good as the candidate models.
One use case for LLM judges is to improve task performance. A potential application is to train on answers chosen by the judge Yuan et al. (2024).
Setup.
Therefore, we test how often the performance of the answers chosen by the judge is better than the performance of the individual models. Formally, for all pairs of models $M_{A},M_{B}$ and datasets $D$, how often is the observed performance $P(\Delta_{J}=T|A,B,D)$ larger than $\max\{P(A=T|D),P(B=T|D)\}$? In Figure 3 the green bar tests all model pairs, and the blue bar only pairs where the judge is at least as good as the candidate models, i.e., $P(J|D)\geq\max\{P(A|D),P(B|D)\}$. The task performances of all models are given in the Appendix B in Table 9.
Results.
We see that only Qwen 2 72B increases the performance reliably. However, it is easier for the judge to improve performance if it compares answers of less or equally good candidate models.
Figure 4: Percentage of model pairs $(M_{A},M_{B})$ where a judge picks a better model $M_{A}$ (meaning $P(A=T|D)>P(B=T|D)$), by selecting more answers of $M_{A}$ than from $M_{B}$.
4.4 Does the judge prefer the better model?
Another application of LLM judges is whether they can accurately identify which model performs better for a given task. This is crucial if we want to rank LLMs by their capabilities or if a practitioner wants to decide which model to deploy.
Setup.
To assess this, we evaluate the frequency with which a judge selects the superior model. For a candidate model pair $M_{A},M_{B}\in\mathbf{M}$, always assume they are ordered, such that $P(A=T|D)>P(B=T|D)$. Then, specifically, we determine the proportion for which the judge chooses $M_{A}$ more often than $M_{B}$, or formally, how often is $P(\Delta_{J}=M_{A}|A,B,D)>P(\Delta_{J}=M_{B}|A,B,D)$ for all candidate pairs and datasets.
Results.
The judges consistently perform well in the selection of the better model. Notably, we find that Qwen 2 72B can only not rank the pair Mistral 7B v0.1 and v0.3 on the MATH dataset. This issue appears minor, as both models exhibit similarly poor performance on the challenging MATH dataset (with accuracies of $6.13\%$ and $3.10\%$, respectively), meaning most judgments compare two wrong answers. Notably, already the worst judge, Mixtral 8x7B, performs well. In summary, we see that judges are more capable of aggregate-level rankings than instance-level rankings.
(a) Qwen2 72B
(b) LLama2 70B
(c) Yi 1.5 43B
(d) Mixtral 8x7B
Figure 5: Judges’ accuracy vs. performance gap between two candidate models $M_{A}$ and $M_{B}$. Each point represents a subset where $M_{A}$ is correct and $M_{B}$ is incorrect. The color reflects the size of these subsets.
5 Analysis of Subsets
We investigate properties that occur when we use subsets based on the correctness of models or agreement between models.
5.1 Do task performances correlate with judgments?
We consider the subset of highest practical relevance where one candidate model is correct, and one candidate model is incorrect. The goal is to investigate the relationship between candidate model task performance and judgment performance.
Setup.
For all model pairs $M_{A},M_{B}\in\mathbf{M},M_{A}\neq M_{B}$ we analyze subsets where $M_{A}$ is correct, and $M_{B}$ is incorrect. Note that we can always order $M_{A}$ and $M_{B}$ this way. Each plot in Figure 5 shows the relationship between judge performance, $P(\Delta_{J}=T|A=T,B=F)$ (Y-axis) and candidate model performance gap of $M_{A}$ and $M_{B}$, i.e., $P(A=T|D)-P(B=T|D)$ (X-axis). Examples of these subsets and their corresponding performances are in Appendix C in Table 10.
Results.
The analysis reveals a strong correlation (Pearson’s $r^{2}>0.69$) between candidate model performance gap and judgment accuracy. If the performance gap is negative, we consider subsets where larger models are incorrect. Judges favor answers from larger models even when they are incorrect on these subsets. We hypothesize that this bias arises because larger models exhibit a specific writing style, articulating their responses more convincingly, thereby misleading the judges. This finding aligns with previous research identifying self-bias Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024). However, our results indicate that this bias extends more broadly to the inherent quality of the underlying models on reasoning datasets. However, this is not necessarily a critical issue in practice, as the larger model tends to answer correctly more often (as indicated by the color of the points in Figure 5.
5.2 Does judgment quality depend on models’ agreement?
(a) Performance on all comparisons.
(b) Performance on comparisons with one correct and one incorrect answer.
Figure 6: Judge performance by agreement bucket. E.g., bucket $S_{3}$ (X-axis) means that all eight models gave together three different answers. Note that AQUA-RAT is multiple-choice with maximally six answers.
We are interested in whether the level of agreement among models, i.e., how many models give a different answer for a sample, impacts the performance on the respective subset.
Setup.
We define disagreement buckets $S_{j}$, where each bucket contains instances for which exactly $1\leq j\leq 8$ unique answers were given across all models. Formally, we set
$S_{j}=\bigcup_{D}\left\{i\in D\mid|\{M_{A}(i)\mid M_{A}\in\mathbf{M}\}|=j\right\}$
where $M_{A}(i)$ is the answer of model $M_{A}$ for instance $i$. We analyze the results in two contexts: all comparisons, including those where both answers are correct or incorrect (cf. Figure 6(a)), and only instances where exactly one answer is correct (cf. Figure 6(b)). We average the performances of all judges and all candidate pairs. Find per-judge plots in Appendix C in Figure 10.
Results.
Figure 6(a) shows that when all models agree (bucket $S_{1}$), the performance is nearly 100%, indicating unanimous agreement usually means correctness. As disagreement increases, performance expectedly decreases. Thus, model agreement is a proxy for sample difficulty. In 6(b), where a correct and an incorrect answer exists, performance remains relatively stable across disagreement buckets for datasets with free-form answers, such as GSM8K and MATH. However, for AQUA-RAT, performance degrades as disagreement rises.
[…]
↓ Features \  Condition → $A,B$ $A\neq B$$\{T,F\}=\{A,B\}$
(1) $P(J),P(A),P(B)$ 97.50 90.20 59.20
(2) $P(A=B)$, 76.00 54.90 49.90
$P(J=A|A\neq B)$
Table 2: Coeffictions of Determination ($R^{2}$, higher is better) for linear regression using the different feature sets as covariates (rows) and different target variables defined by the condition (columns). All values are significant ($p<0.001$) as per an Overall-F-Test.
6 Prediction of Judgements
We investigate whether predicting the judgments’ outcomes is feasible. Firstly, we aim to predict performance statistics. Secondly, we aim to predict individual judgments.
6.1 Can we predict judgment performance?
On the subset where exactly one answer is correct, we found a strong correlation between judgment performance and candidate task performances. This hints at regularities within the judging process, thus we aim to predict judge performance using model statistics.
Setup.
We fit six different linear regression models using the judgment performances as the target variables $Y$, including all variations of judges, model pairs $M_{A},M_{B}\in\mathbf{M}$, and datasets $D$. Regarding the covariates $\mathbf{X}$ in the model, we distinguish between two setups: In Case (1), we solely use the task performances $P(X|D),X\in\{J,A,B\}$ of judge and candidate models, to predict judgment performance. In Case (2), we utilize statistics available without knowledge of the ground truth. The features for this case are the probability of agreement between the candidate models $P(A=B|D)$ and the probability of model $M_{A}$ being chosen. Since we are not specifically interested in the individual features’ effects, but rather in their ability to explain the variation of judgment performance, we rely on the coefficient of determination, $R^{2}$, for evaluation (Fahrmeir et al., 2013, , see Appendix E).
Results.
The results are shown in Table 2 (excluding data sets from the probability formulas for simplicity). We observe that the performance-related features of the models can almost perfectly explain the variation in final judgment performance ($R^{2}=97.50\%$), also when conditioning only on the subset of differing answers ($R^{2}=90.20\%$). Logically, $P(A)$ and $P(B)$, i.e., $P(A|D),P(B|D)$ respectively, have significant [^3] We test statistical significance using an Overall-F-Test for each fitted model. Further details are in Appendix E. explanatory power for judgment performance, as they encompass all correct answers. In Case (2), we still observe a relatively high $R^{2}$ value, indicating that the features can explain 50% of the target’s variance.
[…]
(1) TF-IDF + RF 60.78 61.37 60.77 58.69
(2) RoBERTa 68.14 66.49 67.03 63.91
Table 3: Accuracy of predicting LLM judges’ decisions using Random Forest (RF) and RoBERTa classifiers.
6.2 Can we predict which individual judgments?
We hypothesize that judgments are biased towards larger or better models because they incorporate linguistic cues or writing style into their judgments rather than purely relying on reasoning assessment. Therefore, we train a classifier to understand whether we can predict individual judgments.
Setup.
We separate all comparisons made per judge into training, validation, and test splits and train two classifiers. The test accuracy is reported in Table 3. The first model utilizes TF-IDF vectorization. We create two independent vectorizers for both answers. The resulting features are concatenated. A RandomForest classifier Breiman (2001) is then trained on these combined features. The second model is a RoBERTa model Liu et al. (2020) trained on the full prompt presented to the judge. Refer to Appendix D for the training details of both models.
Results.
The random forest model achieves an accuracy of approximately 60%, demonstrating performance above random chance. This suggests that specific keywords or phrases influence judges. The RoBERTa model surpassed this, reaching nearly 70% accuracy. Taken together, these results suggest that judge decision-making is a multi-faceted process. While specific linguistic cues appear to hold influence, a substantial portion of the decision-making process seems to be based on other contextual factors or broader reasoning.
7 Perturbation of Results
Swapped Masked
= $\neq$ Refused = $\neq$ Refused
Qwen2-72B 75.75 12.20 12.05 56.40 26.06 17.53
Llama-3-70B 78.81 13.72 7.47 63.44 29.22 7.33
Yi-1.5-34B 74.80 14.47 10.73 44.89 37.31 17.80
Mixtral-8x7B 71.19 20.69 8.12 60.34 29.26 10.40
Table 4: Analysis of judgments where results in candidate answers were either swapped or numbers masked. We report how many judgments stay the same (=)(=)( = ), different ()($\neq$)( ≠ ), or where judges refused to follow the output format (Refused).
We aim to gain a deeper understanding of the extent to which writing style affects the final judgment. Therefore, we create an experiment perturbing the candidate answers and examine whether this changes the judgment.
Setup.
We examine two perturbtations: Swap and Mask. In the Swap experiment, we swap the final answer from model $M_{A}$ with that of model $M_{B}$, while keeping their CoT reasoning unchanged. In the Mask experiment, we anonymize all numbers in both the CoT reasoning and the final answer by replacing them with “X”. [^4] In preliminary runs, we observed that masking caused significant confusion for the judge models. To address this, we adapt the judgment prompt in this setting to include the instruction: “Only analyze the reasoning! All numbers have been replaced with ’X’ to help you focus on the reasoning.” Table 4 shows the frequency with which the judge selects the same answer ($=$), a different answer ($\neq$), or fails/refuses to follow the output format and make a decision (Refused). Refer to Appendix F for specific examples.
Results.
We observe that the new judgments in more than half the cases agree with the original judgment. In the Swap experiment, they even agree on average by 75% of the cases. We deduce that the judge is largely unaffected by the artificially introduced noise and heavily bases its decision on the writing style. Interestingly, in a substantial amount of samples (up to $17\%$) the judge refuses to make a judgment. On a positive note, manual inspection revealed that the model often realizes that the original answers were perturbed.
8 Discussion
Style and Quality.
Our experiments suggest a relation between judgment and candidate task performance (cf. Section 5) and a relation between judgment and writing style (cf. Sec. 6 and 7). We hypothesize these two are interconnected and facets of the same underlying bias. When models become better, e.g., by being trained on larger amounts of data, their ability to write convincingly increases. Conversely, when an LLM demonstrates an increased ability to write convincingly, it likely acquires a more nuanced grasp of what humans perceive as compelling. This enhanced understanding likely also extends to task performance.
Generalizability of approach.
Our in-depth analysis utilizes Formula (1) to segment judgment data based on correctness criteria, allowing for targeted investigation of specific subsets. This approach is generalizable and transferable to other NLP tasks, such as summarization. By incorporating discrete signals such as text topics, a similar derivation of the judgment probability is possible.
9 Conclusion
We conducted a thorough analysis of LLM judges on mathematical reasoning tasks. We include a detailed judgment performance evaluation of eight models on three datasets. We find that larger models are generally better than smaller models and that judges succeed in detecting the more capable model. Our analysis reveals a strong correlation between judgment performance and task performance of the models providing candidate answers which shows that judges tend to choose larger or better models. We hypothesize that LLM judges incorporate writing style into their judgments instead of purely analyzing the reasoning. We provide two experiments to provide evidence for this hypothesis.
Finally, we want to emphasize the importance of impartiality and fairness in the role of LLM judges, similar to human judges in the real world. Our research introduces methods to quantify biases in favor of larger or better models, thereby offering a means to measure the reduction of such biases.
Limitations
Our analysis is primarily focused on mathematical reasoning datasets, which allows us to explore judgments through the lens of correctness within specific subsets. While this approach provides valuable insights, it limits the generalizability of our findings to other tasks or domains. Based on the fact that the investigated datasets are complex, in the sense that they need multi-step reasoning to be solved, and based on the fact that there is no thorough investigation of LLM judges on mathematical reasoning datasets yet, we think this work is a valuable contribution.
In our experiments, we focus on testing a single, specific prompt. It is common knowledge that LLMs are highly sensitive to variations in prompt phrasing, which can substantially influence their performance. Nevertheless, it is impossible for us to meet the computational demands necessary to run our experiments with multiple prompts.
In this study, we intentionally concentrate on open-weight models, motivated by our strong belief in the principles of open science. Open-weight models offer transparency and reproducibility, which are critical for advancing scientific understanding. However, we note that it is also interesting to study closed models to understand potential differences. Still, we are committed to research on open-weight models because we believe it benefits the community more.
10 Acknowledgements
This research was funded by the WWTF through the project ”Knowledge-infused Deep Learning for Natural Language Processing” (WWTF Vienna Research Group VRG19-008) and through the project "Transparent and Explainable Models" (WWTF ICT19-041). MA is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI - grant number 460037581. Further, we thank Jan Philip Wahle and Pedro Henrique Luz de Araujo for fruitful discussions and their constructive feedback.
Appendix A Experimental Setup
We provide further details on the general setup described in Section 3. Specifically, we include statistics and examples of the datasets, additional information on the models used, and the exact prompts employed in this study.
A.1 Datasets
Additional information about the datasets is given in Table 5, which presents an overview of the dataset statistics. Note that for the MATH dataset, we only include the most challenging questions, called levels 4 and 5, in the dataset. Notably, it has ground truth answer sequences that are, on average, almost three times longer than those in other datasets.
In Table 6, we provide examples of questions and their corresponding answers from the ground truth. Note that these examples were used for few-shot prompting.
# questions Avg. Avg.
# question characters # answer characters
AQUA-RAT 254 239.1 203.1
MATH 1516 216.5 643.9
GSM8K 1319 239.9 292.9
Table 5: An overview of dataset size and text length.
Question Answer
AQUA-RAT Two friends plan to walk along a 43-km trail, starting at opposite ends of the trail at the same time. If Friend P’s rate is 15% faster than Friend Q’s, how many kilometers will Friend P have walked when they pass each other? Options: A)21 B)21.5 C)22 D)22.5 E)23 If Q complete x kilometers, then P completes 1.15x kilometers. x + 1.15x = 43 2.15x=43 x = 43/2.15 = 20 Then P will have have walked 1.15*20=23 km. The answer is E. #### E
GSM8K Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Natalia sold 48/2 = <<48/2=24>>24 clips in May. Natalia sold 48+24 = <<48+24=72>>72 clips altogether in April and May. #### 72
MATH Mr. Madoff invests 1000 dollars in a fund that compounds annually at a constant interest rate. After three years, his investment has grown to 1225 dollars. What is the annual interest rate, as a percentage? (Round your answer to the nearest integer.) Let $r$ be the annual interest rate. Then after three years, Mr. Madoff’s investment is $1000\cdot\left(1+\frac{r}{100}\right)^{3}$, so 1000($12251000\cdot\left(1+\frac{r}{100}\right)^{3}=$ 1225 . Then $\left(1+\frac{r}{100}\right)^{3}=1.225,$so [$1+\frac{r}{100}=\sqrt[3]{1.225}=1.069987\dots,$ which means $r=\boxed{7}$, to the nearest integer. #### 7.0
Table 6: Example of ground truth answers used for few-shot prompting.
A.2 Models
We execute all models using the VLLM software for LLM serving Kwon et al. (2023). The weights for all models are accessible through Huggingface Transformers Wolf et al. (2020). Table 7 includes hyperlinks to each model for easy reference.
Table 7: Used models and corresponding hyperlinks.
A.3 Prompts
Initial Prompt { "role": "user", "content": "You are a reasoning assistant. Always answer exactly in the same format. Use ’####’ to separate the final answer (without additional comments) from the reasoning. {{shot 1 question}}" }, { "role": "assistant", "content": "{{shot 1 question}}" } , { "role": "assistant", "content": "{{shot 4 answer}}" }, { "role": "user", "content": "{{question}}" } }
Figure 7: Prompt used to generate initial solutions for all datasets. It includes few-shots and the question of the current sample.
Judge Prompt Question: {{question}} Answer A: {{answer A}} ————– Answer B: {{answer B}} ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘
Figure 8: Prompt used for judgements. The full text above is wrapped in the user role, as all models support this role. No additional system message is used.
We used two different prompts within this project. The prompt shown in Figure 7 is used for the candidate solutions for all datasets. Examples of the few-shots are in Table 6. The prompt for the judges is given in Figure 8. Note that we run experiments for both orders of the answers of the models $M_{A}$ and $M_{B}$.
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
Qwen 2 72B Qwen 2 72B Llama 3 8B 20.00 Question: Charlie wants to sell beeswax candles. For every pound of beeswax, he can make 10 tapered candles. One pound of beeswax and the wicks cost \$10.00 in supplies. If he sells each candle for \$2.00 each, what is his net profit if he makes and sells 20 candles? Answer A: He needs 2 pounds of beeswax to make 20 candles because 20 / 10 = <<20/10=2>>2 It will cost \$20.00 in supplies for the beeswax and wicks because 2 x 10 = <<2*10=20>>20 His total revenue from selling the candles will be \$40.00 because 20 x 2 = <<20*2=40>>40 His net profit will be \$20.00 because 40 - 20 = <<40-$20 ###$ ————– Answer B: It takes 1 pound of beeswax and the wicks to make 10 candles, so it takes 1 pound of beeswax and the wicks to make 20 candles. It costs 10 dollars for 1 pound of beeswax and the wicks, so it costs 10 dollars for 20 candles. Charlie sells each candle for 2 dollars, so he makes 20 * 2 = 40 dollars by selling the 20 candles. So, his net profit is 40 - 10 = 30 dollars. #### 30 ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Answer A and Answer B differ in their calculation of the cost of supplies for making 20 candles. Answer A correctly calculates that Charlie needs 2 pounds of beeswax to make 20 candles and that the cost for the beeswax and wicks for 2 pounds is \$20.00. Answer B incorrectly assumes that the cost for making 20 candles is the same as for 10 candles, which is \$10.00. Therefore, Answer A correctly calculates the net profit as \$20.00, while Answer B incorrectly calculates it as \$30.00. The correct answer is A. “‘json { "answer": "A" } “‘ True
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: There is well of depth 30m and frog is at bottom of the well. He jumps 3m up one day and falls back 2m down the same day. How many days will it take for the frog to come out of the well? Options: A)25 days B)26 days C)27 days D)28 days E)29 days Answer A: The frog is initially at the bottom of the well, which is 30m deep. On the first day, the frog jumps 3m up, so it is 33m from the bottom. Then, it falls back 2m, so it is 31m from the bottom. Each day, the frog jumps 3m up and falls 2m down, making a net gain of 1m. To get out of the 30m deep well, it would take 30 days for the frog to jump and fall 30 times, covering a total distance of 30m. #### A ————– Answer B: If he jumps 3m up and slips 2m down in a day, he is 1m up in a day in 27 days he would have reached 27m On the 28th day, he jumps up 3m and come out of the well. Answer: D #### D ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s break down both answers: **Answer A:** The explanation is correct that the frog makes a net gain of 1m each day. However, the conclusion that it would take 30 days to get out of the 30m deep well is incorrect. The frog would reach the top of the well when it has covered a total distance of 30m, but it would not take 30 days. **Answer B:** This explanation is correct. The frog makes a net gain of 1m each day, so in 27 days, it would have covered 27m. On the 28th day, it would jump up 3m and come out of the well. Therefore, the correct answer is **D) 28 days**. Here is the conclusion in JSON format: “‘json { "answer": "B" } “‘ True
Table 8: Example of ground truth answers used for few-shot prompting.
A.4 Infrastructure
The experiments were run on NVIDIA A100 and NVIDIA H100. The judgments used in Section 4 and Section 5 took around 5 days equivalents on 4 A100 40GB. Using 2 H100 90GB and 4 A100 40GB it took around 2.5 days. For the perturbation experiments in Section 7, it took around the same amount of time.
Appendix B General performance
This section provides additional information related to Section 4. Specifically, we present the task performance of all models across all datasets, as well as the judging performance of all models when used as judges.
B.1 Task Performance
In various contexts in this work, the task performance of the individual models is essential. Therefore, we provide the accuracy of all models and all datasets in Table 9.
AQUA-RAT GSM8K MATH
Qwen 2 72B 76.38 92.04 51.19
Llama 3 70B 73.62 91.05 34.37
Yi 1.5 34B 64.96 78.47 27.04
Mixtral 8x7B 47.24 61.18 13.79
Llama 3 8B 51.18 73.01 15.04
Gemma 1.1 7B 42.91 50.72 12.60
Mistral 7B v0.3 38.19 42.76 6.13
Mistral 7B v0.1 21.65 26.08 3.10
Table 9: Task performance of all models using the prompt in Figure 7.
B.2 Judging performance per model pair
(a) Qwen 2 70B
(b) LLama3 70B
(c) LLama 3 8B
(d) Gemma 1.1 7B
(e) Yi 1.5 34B
(f) Mixtral 8x7B
(g) Mistral 7B v0.3
(h) Mistral 7B v0.1
Figure 9: Evaluation of final task performance $P(\Delta_{J}=T|A,B)$ averaged over all datasets for model pairs $(M_{A},M_{B})$ for the judges (a) - (h).
We conduct experiments with all eight models serving as judges. We present the performance metrics of all judges across various model comparisons in Figure 9. As shown in Table 1, only the large models consistently produce judgments that deviate consistently from random chance. These results in Figure 9 support the superior performance of larger models.
Appendix C Additional subset experiments
This section provides additional information for chapter 5.
C.1 Example Subset performance
To better understand the correlation observed in Figure 5, we provide examples of these subsets, which can be seen in Table 10. These examples include the following details: the judge, the compared models, the dataset, the performance of each model on a dataset (denoted by $P(A|D)$), the judgment performance on the subset (denoted by $P(\Delta_{J}=T|A=X,B=Y,D)$), and the relative size of the subset (denoted by $P(A=X,B=Y|D)$). We provide the five subsets with the highest performance, the five subsets with the five lowest performance, and five random subsets.
Judge model A model B dataset X Y $P(A|D)$ $P(B|D)$ $P(\Delta_{J}=T|A=X,B=Y,D)$ $P(A=X,B=Y|D)$
Qwen 2 72B Qwen 2 72B Mistral 7B v0.1 MATH True False 51.2 3.1 99.1 50.2
Qwen 2 72B Yi 1.5 34B Mistral 7B v0.1 MATH True False 27.0 3.1 98.4 27.9
Qwen 2 72B Llama 3 8B Mistral 7B v0.1 MATH True False 15.0 3.1 98.3 16.7
Qwen 2 72B Llama 3 70B Mistral 7B v0.1 MATH True False 34.4 3.1 98.3 35.4
Qwen 2 72B Mixtral 8x7B Mistral 7B v0.1 MATH True False 13.8 3.1 98.2 15.4
Mixtral 8x7B Mixtral 8x7B Mixtral 8x7B GSM8K False True 61.2 61.2 65.8 14.7
Yi 1.5 34B Llama 3 8B Gemma 1.1 7B GSM8K False True 73.0 50.7 64.9 8.0
Yi 1.5 34B Llama 3 70B Mistral 7B v0.1 AQUA-RAT True False 73.6 21.7 89.1 60.8
Qwen 2 72B Gemma 1.1 7B Gemma 1.1 7B GSM8K False True 50.7 50.7 90.0 12.9
Yi 1.5 34B Yi 1.5 34B Mistral 7B v0.1 GSM8K False True 78.5 26.1 52.5 2.6
Qwen 2 72B Llama 3 70B Mistral 7B v0.1 MATH False True 34.4 3.1 13.2 2.2
Qwen 2 72B Yi 1.5 34B Mistral 7B v0.1 AQUA-RAT False True 65.0 21.7 10.0 4.7
Yi 1.5 34B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.5 1.4
Llama 3 70B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.2 1.3
Qwen 2 72B Qwen 2 72B Mistral 7B v0.1 MATH False True 51.2 3.1 6.1 1.4
Table 10: Examples of comparisons; and performance; problem:
C.2 Performance by agreement
(a) Qwen 2 72B: Using all comparisons.
(b) Comparison with a correct and incorrect answer
(c) Llama 3 70B: Using all comparisons.
(d) Comparison with a correct and incorrect answer
(e) Yi 1.5 34B: Using all comparisons.
(f) Comparison with a correct and incorrect answer
(g) Mixtral 8x7B: Using all comparisons.
(h) Comparison with a correct and incorrect answer
Figure 10: Judge performance by agreement bucket, e.g. bucket 3 (X-axis) means that all eight models gave together three different answers.
We provide an extension of the results in the main paper. We put all samples into bins of how many different answers were given by the eight models. For the sake of brevity, we average across all judges. In Figure 10 we show the same graphs for each judge individually. Again, we give the performance on the overall dataset and on the subset where there exists exactly one correct and one incorrect solution. All interpretations and conclusions overlap with the main part of the paper.
Appendix D Prediction of Individual Judgements
Random Forest.
We utilize TF-IDF representations with English stop word removal for the Random Forest model. We further remove all digits from the text. We set the maximum number of features to 10,000 and consider n-grams ranging from unigrams to bigrams (N-gram range: 1-2). The model uses 500 trees (estimators). For training, we use the Scikit-learn Pedregosa et al. (2011) library. The running time was negligible.
RoBERTa.
For the RoBERTa model Liu et al. (2020), we use a batch size of 64 and a learning rate of 2e-5. The weight decay is set to 1e-3, and the model is trained for 8 epochs. The final model is selected based on the best validation performance. The model is trained using the HuggingFace Transformers library Wolf et al. (2020). The total running time was about twelve hours on a single H100 90GB.
Appendix E Statistical Methodology
We describe the statistical background for the tests applied in Section 6. All predictions and statistical tests in Section 6 were performed using the statsmodels library Seabold and Perktold (2010).
E.1 Coefficient of Determination
The coefficient of determination, $R^{2}$, for evaluation of linear regression models (Fahrmeir et al., 2013) is defined as follows:
$R^{2}=\frac{\sum_{i=1}^{n}(\hat{y}_{i}-\bar{y})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$
$R^{2}$ measures the share of the variance in $Y$ explained by its covariation with the features $\mathbf{X}$ included in the model by dividing the variation of the predicted values $\hat{y}_{i}$ by the variation of the true target values $y_{i}$. If the features $\mathbf{X}$ have high explanatory power for $Y$, the $\hat{y}_{i}$ will be close to the $y_{i}$ and $R^{2}$ will be close to 1, while in the extreme case of no correlation between $\mathbf{X}$ and $Y$ the arithmetic mean is the best estimate (i.e., $\hat{y}_{i}=\bar{y}\;\forall\;i=1,\dots,n$) resulting in $R^{2}=0$.
E.2 Overall-F-Test
The Overall-F-Test is built upon $R^{2}$ and tests whether the overall model is of any significant value for explaining the variation of the target variable. The F-distributed test statistic is calculated as
$\frac{R^{2}}{1-R^{2}}\cdot\frac{n-p-1}{p},$
where $R^{2}$ is the coefficient of determination, $n$ is the number of observations, and $p$ is the number of covariates included in the model (i.e., the number of estimated coefficients excluding the intercept). The hypotheses that can be tested this way are
$H_{0}:\beta_{1}=\beta_{2}=\dots=\beta_{p}=0$
vs.
$\qquad H_{1}:\beta_{j}\neq 0\;\mbox{for at least one}\;j\in\{1,\dots,p\}.$
So from a rejection of $H_{0}$, it can be concluded that at least one of the included features exhibits explanatory power for the variation of the target variable.
E.3 Multiple Testing
Since we conduct multiple statistical tests within the scope of one research project, it is important to consider multiple testing as a potential problem resulting in false positive findings. The p-values from our tests, however, also satisfy a significance level resulting from a Bonferroni Correction of the typical significance level of 5%.
Appendix F Peturbation of Results
In Section 7 we performed two experiments, where we 1) swap the results of the individual answers and 2) mask all numbers using an ’X’ to understand whether the model only focuses on the writing style. You can find examples, including the prompt and the response of the judge for experiment 1) in Table 11 and examples for experiment 2) in Table 12
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
Qwen 2 72B Qwen 2 72B Llama 3 8B 163.00 Question: Candice put 80 post-it notes in her purse before she headed out to her job at the coffee shop. On her way, she stopped off at the store and purchased a package of Post-it notes; At work, she placed a single Post-it note on each of 220 different cups of coffee. If she had 23 post-it notes remaining overall, how many Post-it notes were in the package that she purchased? Answer A: Candice had 80 post-it notes initially. She used 80 + 220 = <<80+220=300>>300 of them to note the cups of coffee. She had 23 post-it notes remaining overall. So, she used 300 - 23 = <<300-23=277>>277 of the 300 post-it notes initially. The remaining post-it notes are the difference between the total and the initial count, which is 300 - 80 = <<300-80=220>>220. She must have purchased 220 - 23 = <<220-23=163>>197 post-it notes. #### 163 ————– Answer B: Candice started with 80 Post-it notes and placed one on each of the 220 cups of coffee, so she used 220 notes. If she had 23 notes remaining, then she must have had 220 + 23 = 243 notes in total. Since she originally had 80 notes, the package she purchased must have contained 243 - 80 = 197 notes. Therefore, the package contained 197 Post-it notes. #### 197 ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Answer A is incorrect because it incorrectly calculates the number of post-it notes Candice used and purchased. The correct answer is B because it accurately calculates the total number of post-it notes Candice had after using them and then subtracts the initial number of post-it notes she had to find out how many were in the package she purchased. “‘json { "answer": "B" } “‘ True
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: In a chocolate store, all chocolates are either vanilla or cocoa flavored only. 10% of the chocolates are cocoa flavored, 90% of the rest are squashed. What percentage of the chocolates are both vanilla flavored and not squashed? Options: A)1% B)2% C)5% D)9% E)10% Answer A: If 10% of the chocolates are cocoa flavored, then the remaining 90% are vanilla flavored. 90% of the vanilla chocolates are squashed. This means that 10% of the vanilla chocolates are not squashed. The question asks for the percentage of chocolates that are both vanilla flavored and not squashed. Since 10% of the vanilla chocolates are not squashed, the answer is 10%. #### D ————– Answer B: 90% of the chocolates are not cocoa flavored (they are vanilla) and 10% of these (90% *10% = 9%) are not squashed. So 9% of the chocolates are both vanilla flavored and not squashed. The answer is E. #### E ————– Compare both answers in detail and choose the answer which correctly answers the question. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s break down both answers: **Answer A:** If 10% of the chocolates are cocoa flavored, then the remaining 90% are vanilla flavored. 90% of the vanilla chocolates are squashed. This means that 10% of the vanilla chocolates are not squashed. The question asks for the percentage of chocolates that are both vanilla flavored and not squashed. Since 10% of the vanilla chocolates are not squashed, the answer is 10%. **Issue with Answer A:** The calculation "10% of the vanilla chocolates are not squashed" is incorrect. 90% of the vanilla chocolates are squashed, which means 10% of 90% (not 10% of the total) are not squashed. **Answer B:** 90% of the chocolates are not cocoa flavored (they are vanilla) and 10% of these (90% * 10% = 9%) are not squashed. So 9% of the chocolates are both vanilla flavored and not squashed. **Correctness of Answer B:** This answer correctly calculates the percentage of chocolates that are both vanilla flavored and not squashed. Therefore, the correct answer is **B**. Here is the conclusion in JSON format: “‘json { "answer": "B" } “‘ True
Table 11: Examples of comparisons where the final results were swapped.
Judge model $M_{A}$ model $M_{B}$ Label Prompt Response Correct
[…]
Llama 3 70B Qwen 2 72B Llama 3 8B D Question: There is well of depth 30m and frog is at bottom of the well. He jumps 3m up one day and falls back 2m down the same day. How many days will it take for the frog to come out of the well? Options: A)25 days B)26 days C)27 days D)28 days E)29 days Answer A: The frog is initially at the bottom of the well, which is Xm deep. On the first day, the frog jumps Xm up, so it is Xm from the bottom. Then, it falls back Xm, so it is Xm from the bottom. Each day, the frog jumps Xm up and falls Xm down, making a net gain of Xm. To get out of the Xm deep well, it would take X days for the frog to jump and fall X times, covering a total distance of Xm. #### A ————– Answer B: If he jumps Xm up and slips Xm down in a day, he is Xm up in a day in X days he would have reached Xm On the Xth day, he jumps up Xm and come out of the well. Answer: D #### D ————– Compare both answers in detail and choose the answer which correctly answers the question. Only analyze the reasoning! Therefore we exchanged all numbers with ’X’ so you can focus on the reasoning. Conclude with a JSON in Markdown format indicating your choice between answer A or B: “‘json { "answer": "B" or "A" } “‘ Let’s analyze the reasoning of both answers: **Answer A:** The answer assumes that the frog makes a net gain of Xm (3m - 2m = 1m) each day. However, it doesn’t consider the fact that the frog needs to reach the top of the well to come out. The answer simply states that it would take X days to cover a total distance of Xm, without considering the remaining distance to the top of the well. **Answer B:** The answer correctly assumes that the frog makes a net gain of Xm (3m - 2m = 1m) each day. It then concludes that in X days, the frog would have reached Xm, and on the Xth day, it would jump up Xm and come out of the well. This answer considers the remaining distance to the top of the well and correctly concludes that the frog would come out on the Xth day. Based on the analysis, I choose: “‘json { "answer": "B" } “‘ True
Table 12: Example of comparisons where numbers were masked using an ’X’.